OpenStudy (anonymous):

(Tan x +1)/ sin x= (1+ cot x) / cos x prove

OpenStudy (anonymous):

\[\frac{ tanx + 1 }{ sinx } =\frac{ 1+\cot x }{ \cos x }\] now we take left hand side LHS\[\frac{ tanx+1 }{ sinx }\] as we know tan x =\[\frac{ 1 }{ cotx }\] \[\frac{ \frac{ 1 }{ cotx }+1 }{ sinx }\] \[\frac{ \frac{ 1+cotx }{ \cot x } }{ sinx }\]\[\frac{ 1+cotx }{ cotx .sinx }\] as cot x =cos x/sinx put above